OpenStudy (anonymous):

Find a third-degree polynomial function such that f(0) = 18 and whose zeros are -1, 2, and 3. Using complete sentences, explain how you found it

OpenStudy (anonymous):

This can be solved using a system of equations.\[f(x)=a \times x^{3} + b \times x^{2} + c \times x + d\]\[f(0)=18=a \times 0 + b \times 0 + c \times 0 + d\]\[f(-1)=0=a \times (-1)^{3} + b \times (-1)^{2} + c \times -1 + d\]\[f(2)=0=a \times 2^{3} + b \times 2^{2} + c \times 2 + d\]\[f(3)=0=a \times 3^{3} + b \times 3^{2}+ c \times 3 + d\]So then your system is\[\left[\begin{matrix}0 & 0 & 0 & 1 & 18 \\ -1 & 1 & -1 & 1 & 0 \\ 8 & 4 & 2 & 1 & 0 \\ 27 & 9 & 3 & 1 & 0\end{matrix}\right]\]So now you can use any method you prefer to solve this system.